OpenStudy (anonymous):

Find the value of m that makes the two vectors parallel and perpendicular <2m,10>, <-3,6>

OpenStudy (zzr0ck3r):

they are perp if the dot product is 0 so 2m(-3)+10*6 = 0 solve for m

OpenStudy (zzr0ck3r):

there is a similar property with cross product and parallel...

OpenStudy (anonymous):

thank you! it makes more sense that way.

OpenStudy (anonymous):

what do I make it equal to to find parallel?

OpenStudy (anonymous):

for parllel slope will be eqaul -(10)/2m = -(6/-3) -5/2 = m m = -2.5